OpenStudy (anonymous):

Let w(x) = sqrt(2x-4). This function can be written as a composition: w(x) = f(g(x)) where f(x) = sort(x). a. What is the domain of f(x)? b. For the composition f(g(x)) to make sense, the output of g(x) must be inside the domain of f(x). For what values of x is this true? (You should set up an inequality and solve for x to answer this)

OpenStudy (anonymous):

For the first part of the question, I got f(x) = sqrt(x) so the domain is all real numbers or domain: [0, oo)

OpenStudy (anonymous):

Am I on the right track for part b? I have g(x) = 2x+4 = 0, solving for x I get x = 2 and so that is part of my domain

OpenStudy (anonymous):

it says set up an inequality: \(2x+4\geq 0\)

OpenStudy (anonymous):

sry, should be \(2x-4\geq\) 0

OpenStudy (anonymous):

solve this and that will be the domain

OpenStudy (anonymous):

the domain that I got was [2, oo)